OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3)

OpenStudy (anonymous):

@hartnn @radar ?

OpenStudy (anonymous):

inequalities

OpenStudy (anonymous):

-2x+6>= 5-x-3 -2x+6>=-x+2 6-4>= -x+2x 4>=x

OpenStudy (anonymous):

I don't understand how to get to that.

OpenStudy (anonymous):

perform operations same as equality just dont change inequality sign

OpenStudy (radar):